# Exponentially weighted estimands and the exponential family: filtering, prediction and smoothing

### v1

@@ 1 Introduction
This paper provides a new, simple way to carry out filtering, predicting and smoothing. This is based on a discounted version of a convex combination of a log-likelihood and the expected log-likelihood function. The expected log-likelihood term has many familiar interpretations, e.g. shrinkage and mean reversion. The discounting is carried out using an exponential function.
We focus on log-likelihoods based around the canonical exponential family. It has a similar structure to generalized linear regression models, but now for time series. Various properties of the resulting procedures are established and illustrated by simulation, estimation and empirical work.
Section 2 develops the new filter, predictor and smoother based on what we call the weighted estimands. Section 3 looks at the case of the canonical exponential family of models. The resulting filters, predictors and smoothers are derived from the weighting principle. These have to be found by a convex numerical optimization.
@@
Section 5 develops a quasi-likelihood based estimation strategy for the hyperparameters of this time series model structure. This method is extended to allow for a two-step estimator, where an easy to use moment estimator reduces the dimension of the required numerical maximization.
Section 6 provides an empirical illustration based on a seven category Dirichlet model. The section uses the empirical results to provide simulation based evidence for the performance of the hyperparameter estimators.
Section 7 details some conclusions. Section 8 contains a proof.
@@ 2 Principle of weighting
The core of this paper will be based around an exponential version of a weighting principle, [^3] The extension to allow a prior density $\pi(\theta)$ for $\theta\in\Theta$, yielding an exponentially weighted MAP (i.e. mode of a type of posterior), means just adding $\log\pi(\theta)$ to our criterion function and then maximizing that with respect to $\theta$ for each $t$. This is particularly important if $\theta$ is high dimensional.
@@
Let Assumptions 1-2 hold, $\lambda\in[0,1]$ be a discount hyperparameter and $\alpha\in[0,1]$ an anchoring hyperparameter. The exponentially weighted filter at time $t$, using data up to time $t$ and the sequence of frames $\{\log L_{t}(\theta;y_{t})\}_{t=1}^{T}$, is
@@
(a) When $\alpha=1$ the resulting exponentially weighted maximum likelihood estimator (EWMLE) appears often in various literatures often under the label of local modeling or local regression, e.g. Chapter 8 of FanYao(05) focus on non-linear time series and FanHeckmanWand(95) who look at non-linear regression. There the localization is typically carried out using a kernel weighting function — our use of the exponential function is a special case of this. The Gaussian likelihood for an unknown mean yields the celebrated EWMA of Brown(56).
@@
(c) LuxenbergBoyd(24) look at exponentially weighted moving losses, replacing the likelihood with a loss function. Their work has some overlap with our approach, but focuses on penalization (which we do not), and does not consider the $\mathbb{E}[\log L_{j}(\theta,Y_{j})]$ terms.
(d) When $\alpha<1$ we do not know of a literature which uses this approach, including the works cited in parts (a)–(c). Of course, broadly, the expected log-likelihood can be thought as a form of shrinkage. Hence it relates to the vast literatures on empirical Bayes, ridge and Lasso regression, etc (e.g. EfronMorris(77), Efron(12), HoerlKennard(70a); HoerlKennard(70b), Tibshirani(96) and Hansen(1616shrink)).
(e) There is a large general Bayesian literature which raises the likelihood to a power less than one to obtain some “robustness” properties for the posterior. Some of the corresponding literature is discussed in HolmesWalker(17).
@@
(g) In applications with naturally constrained parameters (e.g. $\sigma>0$), we work with a smooth one-to-one reparametrization such as a log or logit transform so that the resulting parameter vector $\theta$ lives in $\mathbb{R}^{d}$. Since the exponentially weighted estimands are reparametrization invariant, this involves no loss of generality and allows us to take $\Theta=\mathbb{R}^{d}$. Together with the upper semicontinuity of the log-frame and the coercivity and strict concavity of the expected log-frame in Assumption 2, this guarantees that all $Q_{\lambda,t}$, $Q_{\lambda,t|t-s}$ and $Q_{\lambda,t|T}$ are coercive, concave and upper semicontinuous, and hence that their maximizers exist and are unique.
@@ 3.1 Canonical exponential family
where $\psi(\theta)$ is infinitely continuously differentiable and strictly convex with respect to $\theta$ while $b(y)>0$ for all $y\in\mathcal{Y}$, (e.g. BarndorffNielsen(78info), McCullaghNelder(89) and Efron(22)).
In the canonical exponential family the cumulant function $\log\mathbb{E}[e^{a^{\tt T}h(Y)};\theta]=\psi(\theta+a)-\psi(\theta)$, so
$\psi^{\prime}(\theta)$ $=\frac{\partial\psi(\theta)}{\partial\theta}$
$=\mathbb{E}[h(Y);\theta]$
$=\mu,$
@@
$\psi^{\prime\prime}(\theta)$ $=\frac{\partial^{2}\psi(\theta)}{\partial\theta\partial\theta^{\tt T}}$
$={\mathrm{V}}(h(Y);\theta),$
the corresponding finite variance. The minimal canonical exponential family requires $\psi(\theta)$ to be strictly convex so ${\mathrm{V}}(h(Y);\theta)$ is positive definite under the model. It is helpful to go from $\mu=\mathbb{E}[h(Y)]$ to $\theta$ through the inverse function
$\psi^{\prime-1}(\mu)=\theta.$
This inverse has a unique solution (as $\psi$ is strictly convex), but in some cases it has to be solved numerically — a classic task in statistics as this is isomorphic to computing the MLE of $\theta$ in the canonical exponential family. Famously, it can be carried out reliably and rapidly using a Newton-Raphson algorithm, which converges quadratically close to the solution. We state this formally as Algorithm 1 for reference later.
@@
There is a substantial literature on time series models based around the exponential family. Examples include ZegerQaqish(88), Li(94), BenjaminRigbyStasinopoulos(03) and the review by Davis03072021. The dynamic conditional score (DCS) filter and the generalized autoregressive score (GAS) filter of Harvey(13) and CrealKoopmanLucas(13) can be applied in this context to deliver recursive filters. Our predictive version is closest to the generalized ARMA work of BenjaminRigbyStasinopoulos(03), which is built around the exponential family and generalized linear models.
@@ 3.2 Filtering, prediction and smoothing
In what follows throughout we take the frame $\log L_{t}(\theta;y)=\log f_{t}(y;\theta)$ as coming from a member of the canonical exponential family ${\tt CEF}(\theta,h_{t},\psi_{t})$. Then, for each $t$, (ignoring the implied $b_{t}(y)$ term as it has no impact on $\theta$ and dropping it means we do not need to make an assumption that $\mathbb{E}[\log b_{t}(Y_{t})]$ exists)
@@
Assume a minimal canonical exponential family frame ${\tt CEF}(\theta,h_{t},\psi_{t})$ for each $t$. Then the exponentially weighted filter ($\widetilde{\theta}_{t}$), predictor ($\widetilde{\theta}_{t|t-s}$ with $s\geq 1$) and smoother ($\widetilde{\theta}_{t|T}$) solve:
@@
Proof. Given in the Appendix.
@@ 3.3 A ${\tt CEF}(\theta,h_{t},\psi_{t})$ data generating process
The use of the frame to derive a filter, which is then used as an input into a data generating process echos the DCS/GAS models of Harvey(13) and CrealKoopmanLucas(13).
@@ 4 Special case with analytic solution: $\psi_{t}=n_{t}\psi$
The following special case has an analytic solution: we will focus on it in the rest of this paper.
@@
Use a minimal canonical exponential family frame ${\tt CEF}(\theta,h_{t},n_{t}\psi)$ for each $t$ where $n_{t}$ is a non-stochastic scalar. Then the exponentially weighted filter ($\widetilde{\theta}_{t}$), predictor ($\widetilde{\theta}_{t|t-s}$ with $s\geq 1$) and smoother ($\widetilde{\theta}_{t|T}$) are:
@@
(a) In the special case of a stable frame canonical exponential family and strictly stationary process then
@@
(c) If $\lambda\in(0,1)$, and $n_{t}=1$ for all $t$, then
$n_{\lambda,t}\rightarrow 1/(1-\lambda)$
as $t\rightarrow\infty.$ If $\lambda=1$, then $n_{\lambda,t}=t$, while if $\lambda=0$, then $n_{\lambda,t}=1$. Likewise $n_{\lambda,t|T}\rightarrow(1+\lambda)/(1-\lambda)$ as $t\rightarrow\infty.$ If $\lambda=1$, then $n_{\lambda,t|T}=T$, while if $\lambda=0$, then $n_{\lambda,t|T}=1$.
@@
The $\widetilde{\mu}_{t|t-1}$ then relates to the generalized ARMA model of BenjaminRigbyStasinopoulos(03) who model a link function of the data as being linear in past conditional means and past link functions of the data.
@@
(f) The $n_{\lambda,t}$ and $x_{\lambda,t}$ are free of $\alpha$ and data, while $h_{\lambda,t}$ is free of $\alpha$ (likewise for the smoothed versions).
@@ 4.1 Simulating ten CEF examples: design
We illustrate the exponentially weighted predictor on ten examples from the canonical exponential family. Given $\widetilde{\mu}_{t|t-1}$, each yields a unique solution for $\widetilde{\theta}_{t|t-1}=(\psi^{\prime})^{-1}(\widetilde{\mu}_{t|t-1})$. In cases 1-6 and 9, this can be solved analytically; in cases 7, 8, and 10, numerically via Algorithm 1. In cases 1-6, the canonical parameter $\theta$ is a scalar.
@@
5. Gaussian (zero mean). For $Y_{t}\in\mathbb{R}$ with known zero mean, the canonical parameterization is
@@
8. Dirichlet. For the $Y_{t}\in\Delta_{d-1}$, a d1d-1 dimensional standard simplex, the canonical parameterization is
@@
The resulting predictor $\widetilde{\sigma}^{2}_{t|t-1}=\widetilde{\mu}_{2,t|t-1}-\widetilde{\mu}^{2}_{1,t|t-1}\geq 0$ by invariance and then Jensen’s inequality.
@@
where $I_{0}$ is the modified Bessel function of the 1st kind and $I_{1}$ is the modified Bessel function of the 1st kind of order 1 (e.g. AbramowitzStegun(70)). Given $\mu$, again the corresponding $\theta=(\psi^{\prime})^{-1}(\mu)$ is unique and can found numerically via Algorithm 1.
@@ 4.2 Simulating ten CEF examples: results
Figure 1: Simulation from $Y_{t}|Y_{1:t-1}\sim\texttt{CEF}(\widetilde{\theta}_{t|t-1},h,\psi)$, where $\widetilde{\theta}_{t|t-1}$ follows the exponentially weighted predictor for time t=5,,T=2000t=5,...,T=2000 with discount parameter $\lambda=0.93$. Top row has the anchoring parameter $\alpha=0.7$, bottom row has $\alpha=0.95$. Columns show the evolution of $\widetilde{{\mathrm{E}}[Y_{t}|Y_{1:t-1}]}=\psi^{\prime}(\widetilde{\theta}_{t|t-1})$ in blue with observations $Y_{t}$ as gray circles.
@@
Figure 2: Simulation from $Y_{t}|Y_{1:t-1}\sim\texttt{CEF}(\widetilde{\theta}_{t|t-1},h,\psi)$, where $\widetilde{\theta}_{t|t-1}$ follows the exponentially weighted predictor for time t=5,,T=2000t=5,...,T=2000 with discount parameter $\lambda=0.93$. The top row has anchoring parameter $\alpha=0.7$, and the bottom row has $\alpha=0.95$. Gray circles show observations $Y_{t}$. Blue lines show the conditional expectation: $\widetilde{\mathbb{E}[Y_{t}|Y_{1:t-1}]}=\psi^{\prime}(\widetilde{\theta}_{t|t-1})$ for the exponential, and $\widetilde{\mathbb{E}[Y_{t}|Y_{1:t-1}]}=\frac{\widetilde{\theta}_{t|t-1}}{\widetilde{\theta}_{t|t-1}+1}$ for the Pareto (y-axis on log scale) when $\widetilde{\theta}_{t|t-1}<-1$ and $\infty$ otherwise. The red line for the Gaussian (zero mean) shows the conditional standard deviation $\widetilde{\sigma}_{t|t-1}=\sqrt{\psi^{\prime}(\widetilde{\theta}_{t|t-1})}$.
@@
Figure 3: Simulation from $Y_{t}|Y_{1:t-1}\sim{\tt CEF}(\widetilde{\theta}_{t|t-1},h,\psi)$, where $\widetilde{\theta}_{t|t-1}$ follows the exponentially weighted predictor for time t=5,,T=2000t=5,...,T=2000 with discount parameter $\lambda=0.93$. The top row has anchoring parameter $\alpha=0.7$, and the bottom row has $\alpha=0.95$. Columns show the evolution of the conditional expectation $\widetilde{\mathbb{E}[Y_{t}|Y_{1:t-1}]}$ (blue line) with observations $Y_{t}$ (gray circles) for the Beta, Gaussian (time-varying mean and variance), and von Mises distributions. For the Gaussian distribution, the red line shows the conditional standard deviation $\widetilde{\sigma}_{t|t-1}$, shifted down by 4 units for visualization.

### v2

@@ Abstract
aDept. of Economics and Dept. Statistics, Harvard University, Cambridge, MA 02138, USA
bEconometric Institute, Erasmus University Rotterdam, Rotterdam, The Netherlands
c
Tinbergen Institute, The Netherlands
11footnotetext: Corresponding author.
E-mail addresses: donkervanheel@ese.eur.nl (Simon Donker van Heel), shephard@fas.harvard.edu (Neil Shephard).
@@ 1 Introduction
We provide a simple way to carry out filtering, predicting and smoothing. It is based on a discounted version of a convex combination of a log-likelihood and the expected log-likelihood function. The discounting is carried out using an exponential function.
We focus on log-likelihoods based around the canonical exponential family. It has a similar structure to generalized linear regression models, but now for time series. Properties of the resulting procedures are established and illustrated by simulation and empirical work.
Section 2 develops the new filter, predictor and smoother. Section 3 looks at the canonical exponential family case. These are found by a convex numerical optimization.
@@
Section 5 develops a quasi-likelihood based estimation strategy for the hyperparameters of this structure. This method is extended to allow for a two-step estimator, where an easy to use moment estimator reduces the dimension of the required numerical maximization.
Section 6 provides an empirical illustration based on a seven category Dirichlet model. The section uses the empirical results to provide simulation based evidence for the performance of the hyperparameter estimators. Section 7 details some conclusions.
@@ 2 Principle of weighting
The paper will be based around an exponential version of a weighting principle, [^2] The extension to allow a prior density $\pi(\theta)$ for $\theta\in\Theta$, yielding an exponentially weighted MAP (i.e. mode of a type of posterior), means just adding $\log\pi(\theta)$ to our criterion function and then maximizing that with respect to $\theta$ for each $t$. This is particularly important if $\theta$ is high dimensional.
@@
Let Assumptions 1-2 hold, $\lambda\in[0,1]$ be a discount parameter and $\alpha\in[0,1]$ an anchoring parameter. The exponentially weighted filter at time $t$, using data up to time $t$ and the sequence of frames $\{\log L_{t}(\theta;y_{t})\}_{t=1}^{T}$, is
@@
(a) When $\alpha=1$ the resulting exponentially weighted maximum likelihood estimator appears often in various literatures often under the label of local modeling or local regression, e.g. Chapter 8 of Fan and Yao (2005) and Fan et al. (1995). There the localization is typically carried out using a kernel weighting function — our use of the exponential function is a special case of this. The Gaussian likelihood for an unknown mean yields the celebrated EWMA of Brown (1956).
@@
(c) Luxenberg and Boyd (2024) look at exponentially weighted moving losses, replacing the likelihood with a loss function. Their work has some overlap with our approach, but focuses on penalization (which we do not), and does not consider the $\mathbb{E}[\log L_{j}(\theta,Y_{j})]$ terms.
(d) When $\alpha<1$ we do not know of a literature which uses this approach, including the works cited in parts (a)–(c). Of course, broadly, the expected log-likelihood can be thought as a form of shrinkage. Hence it relates to the vast literatures on empirical Bayes, ridge and Lasso regression, etc (e.g. Efron and Morris (1977), Efron (2012), Hoerl and Kennard (1970), Tibshirani (1996)).
(e) There is a large general Bayesian literature which raises the likelihood to a power less than one to obtain some “robustness” properties for the posterior. Some of the corresponding literature is discussed in Holmes and Walker (2017).
@@
(g) In applications with naturally constrained parameters (e.g. $\sigma>0$), we work with a smooth one-to-one reparametrization such as a log or logit transform so that the resulting parameter vector $\theta$ lives in $\mathbb{R}^{d}$. Since the exponentially weighted estimands are reparametrization invariant, this involves no loss of generality and allows us to take $\Theta=\mathbb{R}^{d}$. Together with the upper semicontinuity of the log-frame and the coercivity and strict concavity of the expected log-frame, this guarantees that all $Q_{\lambda,t}$, $Q_{\lambda,t|t-s}$ and $Q_{\lambda,t|T}$ are coercive, concave and upper semicontinuous, and hence that their maximizers exist and are unique.
@@ 3.1 Canonical exponential family
where $\psi(\theta)$ is infinitely continuously differentiable and strictly convex with respect to $\theta$ while $b(y)>0$ for all $y\in\mathcal{Y}$, (e.g. McCullagh and Nelder (1989) and Efron (2022)).
The cumulant function of the CEF is $\log\mathbb{E}[e^{a^{\tt T}h(Y)};\theta]=\psi(\theta+a)-\psi(\theta)$, so
$\psi^{\prime}(\theta)$ $=\frac{\partial\psi(\theta)}{\partial\theta}=\mathbb{E}[h(Y);\theta]=\mu,$
@@
$\psi^{\prime\prime}(\theta)$ $=\frac{\partial^{2}\psi(\theta)}{\partial\theta\partial\theta^{\tt T}}={\mathrm{V}}(h(Y);\theta),$
the corresponding finite variance. The minimal CEF requires $\psi(\theta)$ to be strictly convex so ${\mathrm{V}}(h(Y);\theta)$ is positive definite under the model. It is helpful to go from $\mu=\mathbb{E}[h(Y)]$ to $\theta$ through the inverse function $\psi^{\prime-1}(\mu)=\theta.$ This inverse has a unique solution (as $\psi$ is strictly convex), but in some cases it has to be solved numerically — a classic task in statistics as this is isomorphic to computing the MLE of $\theta$ in the CEF. Famously, it can be carried out reliably and rapidly using a Newton-Raphson algorithm, which converges quadratically close to the solution. We state this formally as Algorithm 1 for reference later.
@@
There is a substantial literature on time series models based around the exponential family. Examples include Zeger and Qaqish (1988), Li (1994), Benjamin et al. (2003) and the review by Davis et al. (2021). The dynamic conditional score (DCS) filter and the generalized autoregressive score (GAS) filter of Harvey (2013) and Creal et al. (2013) can be applied in this context to deliver recursive filters. Our predictive version is closest to the generalized ARMA work of Benjamin et al. (2003), which is built around the exponential family and generalized linear models.
@@ 3.2 Filtering, prediction and smoothing
In what follows throughout we take the frame $\log L_{t}(\theta;y)=\log f_{t}(y;\theta)$ as coming from ${\tt CEF}(\theta,h_{t},\psi_{t})$. Then, for each $t$, (ignoring the implied $b_{t}(y)$ term as it has no impact on $\theta$ and dropping it means we do not need to make an assumption that $\mathbb{E}[\log b_{t}(Y_{t})]$ exists)
@@
Assume a minimal CEF frame ${\tt CEF}(\theta,h_{t},\psi_{t})$ for each $t$. Then the exponentially weighted filter ($\widetilde{\theta}_{t}$), predictor ($\widetilde{\theta}_{t|t-s}$ with $s\geq 1$) and smoother ($\widetilde{\theta}_{t|T}$) solve:
@@
Proof. Focus on the filtering case. Now
Differentiate $Q_{\lambda,t}(\theta)$ with respect to $\theta$ and solve. This yields the stated result. The same argument applied to prediction and smoothing. The remaining results follow by invariance.
@@ 3.3 A ${\tt CEF}(\theta,h_{t},\psi_{t})$ data generating process
The use of the frame to derive a filter, which is then used as an input into a data generating process echos the DCS/GAS models of Harvey (2013) and Creal et al. (2013).
@@ 4 Special case with analytic solution: $\psi_{t}=n_{t}\psi$
The following special case has an analytic solution. We focus on it in the rest of this paper.
@@
Use a minimal ${\tt CEF}(\theta,h_{t},n_{t}\psi)$ frame for each $t$ where $n_{t}$ is a non-stochastic scalar. Then the exponentially weighted filter ($\widetilde{\theta}_{t}$), predictor ($\widetilde{\theta}_{t|t-s}$ with $s\geq 1$) and smoother ($\widetilde{\theta}_{t|T}$) are:
@@
(a) In the special case of a stable frame CEF and strictly stationary process then
@@
(c) If $\lambda\in(0,1)$, and $n_{t}=1$ for all $t$, then $n_{\lambda,t}\rightarrow 1/(1-\lambda)$ as $t\rightarrow\infty.$ If $\lambda=1$, then $n_{\lambda,t}=t$, while if $\lambda=0$, then $n_{\lambda,t}=1$. Likewise $n_{\lambda,t|T}\rightarrow(1+\lambda)/(1-\lambda)$ as $t\rightarrow\infty.$ If $\lambda=1$, then $n_{\lambda,t|T}=T$, while if $\lambda=0$, then $n_{\lambda,t|T}=1$.
@@
The $\widetilde{\mu}_{t|t-1}$ then relates to the generalized ARMA model of Benjamin et al. (2003) who model a link function of the data as being linear in past conditional means and past link functions of the data.
@@
(f) The $n_{\lambda,t}$ and $x_{\lambda,t}$ are free of $\alpha$ and data, while $h_{\lambda,t}$ is free of $\alpha$.
@@ 4.1 Simulating ten CEF examples: design
We illustrate the exponentially weighted predictor on ten examples from the CEF. Given $\widetilde{\mu}_{t|t-1}$, each yields a unique solution for $\widetilde{\theta}_{t|t-1}=(\psi^{\prime})^{-1}(\widetilde{\mu}_{t|t-1})$. In cases 1-6 and 9, this can be solved analytically; in cases 7, 8, and 10, numerically via Algorithm 1. In cases 1-6, the canonical parameter $\theta$ is a scalar.
@@
5. Gaussian (zero mean). For $Y_{t}\in\mathbb{R}$ with known zero mean, then
@@
8. Dirichlet. For the $Y_{t}\in\Delta_{d-1}$, the canonical parameterization is
@@
The resulting predictor $\widetilde{\sigma}^{2}_{t|t-1}=\widetilde{\mu}_{2,t|t-1}-\widetilde{\mu}^{2}_{1,t|t-1}\geq 0$ by invariance and Jensen’s inequality.
@@
where $I_{0}$ is the modified Bessel function of the 1st kind and $I_{1}$ is the modified Bessel function of the 1st kind of order 1 (e.g. Abramowitz,Milton and Stegun (1970)). Given $\mu$, again the corresponding $\theta=(\psi^{\prime})^{-1}(\mu)$ is unique and can found numerically via Algorithm 1.
@@ 4.2 Simulating ten CEF examples: results
Figure 1: Simulation from $Y_{t}|Y_{1:t-1}\sim\texttt{CEF}(\widetilde{\theta}_{t|t-1},h,\psi)$ for time t=5,,T=2000t=5,...,T=2000 with discount parameter $\lambda=0.93$. Top: anchoring parameter $\alpha=0.7$; bottom: $\alpha=0.95$. Columns: evolution of $\widetilde{{\mathrm{E}}[Y_{t}|Y_{1:t-1}]}=\psi^{\prime}(\widetilde{\theta}_{t|t-1})$ in blue with observations $Y_{t}$ as gray circles.
@@
Figure 2: Simulation from $Y_{t}|Y_{1:t-1}\sim\texttt{CEF}(\widetilde{\theta}_{t|t-1},h,\psi)$ for time t=5,,T=2000t=5,...,T=2000 with discount parameter $\lambda=0.93$. Top: anchoring parameter $\alpha=0.7$; bottom: $\alpha=0.95$. Gray circles show observations $Y_{t}$. Blue lines show the conditional expectation: $\widetilde{\mathbb{E}[Y_{t}|Y_{1:t-1}]}=\psi^{\prime}(\widetilde{\theta}_{t|t-1})$ for the exponential, and $\widetilde{\mathbb{E}[Y_{t}|Y_{1:t-1}]}=\frac{\widetilde{\theta}_{t|t-1}}{\widetilde{\theta}_{t|t-1}+1}$ for the Pareto (y-axis on log scale) when $\widetilde{\theta}_{t|t-1}<-1$ and $\infty$ otherwise. The red line for the Gaussian (zero mean) shows the conditional standard deviation $\widetilde{\sigma}_{t|t-1}=\sqrt{\psi^{\prime}(\widetilde{\theta}_{t|t-1})}$.
@@
Figure 3: Simulation from $Y_{t}|Y_{1:t-1}\sim{\tt CEF}(\widetilde{\theta}_{t|t-1},h,\psi)$ for time t=5,,T=2000t=5,...,T=2000 with discount parameter $\lambda=0.93$. Top: anchoring parameter $\alpha=0.7$; bottom: $\alpha=0.95$. Columns: evolution of the conditional expectation $\widetilde{\mathbb{E}[Y_{t}|Y_{1:t-1}]}$ (blue line) with observations $Y_{t}$ (gray circles) for the Beta, Gaussian (time-varying mean & variance), & von Mises distributions. For the Gaussian distribution, the red line shows the conditional standard deviation $\widetilde{\sigma}_{t|t-1}$, shifted down by 4 units for visualization.